OpenStudy (fibonaccichick666):

Another Laplace transform: \[L(h(t))=L(3tu_2(t)+4(1-u_2(t))\]

OpenStudy (fibonaccichick666):

So \(u_2(t)\) is the heavyside function for c=2. I need to figure out how to write F(t-c)

OpenStudy (fibonaccichick666):

@hartnn , @Hero

OpenStudy (fibonaccichick666):

@zzr0ck3r

OpenStudy (fibonaccichick666):

so if I write 3(t-2+2) would that be right? which then equals 3(t-2)+6..?

OpenStudy (zzr0ck3r):

I hate applied math:( i.e. I suck at applied math. It has been almost 3 years since I took my one and only dif eq class. Sorry:(

OpenStudy (zzr0ck3r):

otherwise I would jump on this question

OpenStudy (fibonaccichick666):

:( darn

OpenStudy (fibonaccichick666):

thanks anyways

OpenStudy (zzr0ck3r):

People here love applied math so someone will be able to help; i'm sure.

OpenStudy (fibonaccichick666):

I'm hoping hartnn stops by again

hartnn (hartnn):

not quite familiar with \(u_2 (t)\)

hartnn (hartnn):

ohh \(u(t-2)\) ?

OpenStudy (fibonaccichick666):

it's the heaviside function so that \[u_2(t)=\left\{\frac{1~~~t<2}{0~~~t\ge 2} \right\}\]

OpenStudy (fibonaccichick666):

but so the Laplace transform is only for f(t-c)

OpenStudy (fibonaccichick666):

so I need to change 3t into something equivalent

OpenStudy (fibonaccichick666):

or if you have another way to integrate \[g(x)=\left\{ \frac{3t~~~0\le t<2}{4~~~t\ge2}\right\} \]

OpenStudy (fibonaccichick666):

not integrate... laplace transform... ohh dear...

hartnn (hartnn):

sorry, i got away for few minutes we can use time shifting property here \(\Large f(t-a)u(t-a) <-> e^{-as}F(s)\)

OpenStudy (fibonaccichick666):

but I guess the issue for me is how can I apply the shifting property here. I know I need to shift it twice

hartnn (hartnn):

\(tu(t-2) <-> e^{(-2s)} L[t+2]\)

hartnn (hartnn):

t = (t+2) -2

OpenStudy (fibonaccichick666):

so does that mean that here I would need the laplace transform of 3(t+2)?

OpenStudy (fibonaccichick666):

or 3(t+2-2)?

hartnn (hartnn):

e^{-2s} L [3(t+2)]

OpenStudy (fibonaccichick666):

ok, but I guess why don't I have to subtract 2?

hartnn (hartnn):

i have adjusted 't' to be in the form of F(t-2) t = t-2 ... but i can't just subtract 2, i need to add 2 to compensate for the -2 so, (t+2) -2 and my f(t) became t+2

hartnn (hartnn):

*f(t-2) notF

OpenStudy (fibonaccichick666):

I'm sorry I still don't understand

OpenStudy (fibonaccichick666):

we have f(t)=3t

OpenStudy (fibonaccichick666):

so then if a move it... I'd have f(t-2)=3(t-2) yes?

hartnn (hartnn):

\(3t u(t-2)\) need to bring this in the form of f(t-2)u(t-2) thaths our requirement. got this? based on the requirement i will form f(t)

hartnn (hartnn):

so i need f(t-2) to be 3t

OpenStudy (fibonaccichick666):

yea, your u(t-2) is the same as my \(u_2(t)\) just different notation

hartnn (hartnn):

to get f(t) from f(t-2) i substitute t =t+2

hartnn (hartnn):

f(t+2 -2) = 3 (t+2)

OpenStudy (fibonaccichick666):

ohhhh ok, I was looking at it backwards

hartnn (hartnn):

f(t) = 3(t+2)

OpenStudy (fibonaccichick666):

ok I follow, so now I can find that. How do I shift a constant, do I have to?

hartnn (hartnn):

so now bring \(4(u_2(t))\) in the form of \(g(t-2)u_2(t)\) and tell me what will be g(t)

OpenStudy (fibonaccichick666):

but it is \(4(1-u_2(t))\)

hartnn (hartnn):

4- that Laplace T of 4 is trivial

OpenStudy (fibonaccichick666):

but can we do that?

hartnn (hartnn):

L[4] =4/s

OpenStudy (fibonaccichick666):

since \(1-u_2(t)\) defines a different function, I guess I am just wondering if we are allowed to distribute

OpenStudy (fibonaccichick666):

but I guess it will become the same, ok point won. So now, for \(4u_2(t)\) I need g(t-2)=4 still right?

hartnn (hartnn):

if it were 4 u_2 (1-t) then we cannot distribute 4 and say u_2 (4-4t) this will be wrong but 4(1-u_2 (t)) can definitely be 4-4u_2(t)

hartnn (hartnn):

yes g(t-2) =4 so g(t) = ... ?

OpenStudy (fibonaccichick666):

g(t)=6...?

hartnn (hartnn):

why 6 ? you add 2 to 't'

OpenStudy (fibonaccichick666):

I thought it was 4 before but you said it wasn't

hartnn (hartnn):

i said it wasn't ???? sorry :P g(t-2) =4 it doesn't depend on 't' so g(t) is also 4 that why its called a constant function :)

OpenStudy (fibonaccichick666):

ohh ok

OpenStudy (fibonaccichick666):

I must have misunderstood, my bad

OpenStudy (fibonaccichick666):

ok then I got it from here :) Thank you again

hartnn (hartnn):

or you can look it like this way \(g(t-2)=4t^0 \\ g(t) = 4 (t+2)^0= 4\)

hartnn (hartnn):

and welcome ^_^